OpenStudy (anonymous):

What is the value of y in the equation 6.4x + 2.8y = 44.4, when x = 3?

OpenStudy (anonymous):

substitute x=3 into the equation @Rafyy

OpenStudy (anonymous):

first substitute 3 for x in the equation, which gets you to 6.4*3 + 2.8y = 44.4.

OpenStudy (anonymous):

6.4x + 2.8y = 44.4 Plug in 'x': 6.4(3) + 2.8y = 44.4 Multiply 6.4 * 3

OpenStudy (anonymous):

so first just multiply 6.4 and 3. do 6.4*3

OpenStudy (anonymous):

\[6.4(3)+2.8y=44.4\]

OpenStudy (anonymous):

\[19.2+2.8y=44.4\]

OpenStudy (anonymous):

I got 19.2 @iGreen @DuckDynastyfan923

OpenStudy (anonymous):

good. now subtract 19.2 from both sides. what is 44.4 - 19.2?

OpenStudy (anonymous):

minus 19.2 on both side

OpenStudy (anonymous):

25.2

OpenStudy (anonymous):

19.2 + 2.8y = 44.4 Subtract.

OpenStudy (anonymous):

Yes, so we have: 25.2 = 2.8y DIvide 25.2 / 2.8

OpenStudy (anonymous):

\[2.8y-19.2+19.2=44.4-19.2\]

OpenStudy (anonymous):

now divide both sides by 2.8. what is 25.2/2.8?

OpenStudy (anonymous):

9 !

OpenStudy (anonymous):

now you have x = 9!

OpenStudy (anonymous):

Yes! 9 is our answer!